# Provably Safe Reinforcement Learning using Entropy Regularizer

### v2

@@ 6 Conclusion and Future Works
$\begin{split}&R(K)\\ &=\sum_{k=1}^{K}\ \big(V^{P}_{\pi_{k}}(c,x_{0})-V^{P}_{\pi^{*}}(c,x_{0})\big)\\ &\overset{(a)}{=}\sum_{k=1}^{K}\ \big(V^{P^{\tau}}_{\pi_{k}}(c,x_{0})-V^{{P}^{\tau}}_{\pi^{*}}(c,x_{0})\big)\\ &=\sum_{k=1}^{K}\ \big(V^{P^{\tau}}_{\pi_{k}}(c,x_{0})-V^{\tilde{P}_{k}^{\tau}}_{\pi_{k}}(\tilde{c}_{k},x_{0})+V^{\tilde{P}_{k}^{\tau}}_{\pi_{k}}(\tilde{c}_{k},x_{0})-V^{P^{\tau}}_{\pi^{*}}(c,x_{0})\big)\\ &\overset{(b)}{\leq}\sum_{k=1}^{K}\ \big(V^{P^{\tau}}_{\pi_{k}}(c,x_{0})-V^{\tilde{P}_{k}^{\tau}}_{\pi_{k}}(\tilde{c}_{k},x_{0})\big)\\ &\overset{(c)}{\leq}\underbrace{\sum_{k=1}^{K}\ \mathbb{E}^{x_{0}}_{\pi_{k}}\sum_{t=0}^{T_{max}-1}\left[\left(c(X^{\tau}_{k,t},A^{\tau}_{k,t})-\tilde{c}_{k}(X^{\tau}_{k,t},A^{\tau}_{k,t})\right);P^{\tau}\right]}_{\text{(I)}}\\ &+\underbrace{\sum_{k=1}^{K}\mathbb{E}^{x_{0}}_{\pi_{k}}\sum_{t=0}^{T_{max}-1}\left[\left<\left|\big(P^{\tau}-\tilde{P}^{\tau}_{k}\big)(X^{\tau}_{k,t},A^{\tau}_{k,t})\right|,V^{\tilde{P}_{k}^{\tau}}_{\pi_{k}}(\tilde{c}_{k})\right>;P^{\tau}\right]}_{\text{(II)}}\end{split}$ (24)
@@
We bound term $(I)$ as follows.
$\begin{split}(I)&=\sum_{k=1}^{K}\ \mathbb{E}^{x_{0}}_{\pi_{k}}\sum_{t=0}^{T_{max}-1}\left[\left(c(X^{\tau}_{k,t},A^{\tau}_{k,t})-\tilde{c}_{k}(X^{\tau}_{k,t},A^{\tau}_{k,t})\right);P^{\tau}\right]\\ &\leq\frac{4T_{max}}{p-p^{s}}\sum_{k=1}^{K}\mathbb{E}^{x_{0}}_{\pi_{k}}\left[\sum_{t=0}^{T_{max}-1}\hat{\epsilon}^{\tau}_{k}(X^{\tau}_{k,t},A^{\tau}_{k,t});P^{\tau}\right]\\ &\overset{\overset{(a)}{<}}{\sim}\tilde{\mathcal{O}}\left(\frac{1}{p-p^{s}}\left(|\mathcal{X}|\sqrt{|\mathcal{A}|T^{3}_{max}K}+|\mathcal{X}|^{2}|\mathcal{A}|T^{2}_{max}\right)\right)\end{split}$ (25)
@@
$\begin{split}&(II)\\ &=\sum_{k,t}\mathbb{E}^{x_{0}}_{\pi_{k}}\left[\sum_{y\in\mathcal{X}}\big|\left(P^{\tau}-\tilde{P}^{\tau}_{k}\right)(X^{\tau}_{k,t},A^{\tau}_{k,t},y)\big|V^{\tilde{P}_{k}^{\tau}}_{\pi_{k}}(\tilde{c}_{k},y)\ ;P^{\tau}\right]\\ &\overset{}{=}\sum_{k,t,x,a}\sigma_{k,t}(x,a)\sum_{y\in\mathcal{X}}\big|\left(P^{\tau}-\tilde{P}^{\tau}_{k}\right)(x,a,y)\big|V^{\tilde{P}_{k}^{\tau}}_{\pi_{k}}(\tilde{c}_{k},y)\\ &=\underbrace{\sum_{k,t,x,a}\sigma_{k,t}(x,a)\sum_{y\in\mathcal{X}}\big|\left(P^{\tau}-\tilde{P}^{\tau}_{k}\right)(x,a,y)\big|V^{{P}^{\tau}}_{\pi_{k}}({c},y)}_{(A.1)}\\ &+\underbrace{\sum_{k,t,x,a}\sigma_{k,t}(x,a)\left<\big|\left(P^{\tau}-\tilde{P}^{\tau}_{k}\right)(x,a)\big|,\big(V^{\tilde{P}_{k}^{\tau}}_{\pi_{k}}(\tilde{c}_{k})-V^{{P}^{\tau}}_{\pi_{k}}({c})\big)\right>}_{(A.2)}\end{split}$ (26)
Term $(A.1)$ can be bounded as:
$\begin{split}(A.1)&=\sum_{k,t,x,a}\sigma_{k,t}(x,a)\sum_{y\in\mathcal{X}}\big|\left(P^{\tau}-\tilde{P}^{\tau}_{k}\right)(x,a,y)\big|V^{{P}^{\tau}}_{\pi_{k}}({c},y)\\ &\overset{(a)}{\leq}T_{max}\sum_{k,t,x,a}\sigma_{k,t}(x,a)\sum_{y\in\mathcal{X}}\hat{\epsilon}^{\tau}_{k}(X^{\tau}_{k,t},A^{\tau}_{k,t},y)\\ &\overset{\overset{(b)}{<}}{\sim}\tilde{\mathcal{O}}\left(\sqrt{|\mathcal{X}|^{2}|\mathcal{A}|T^{3}_{max}K}+|\mathcal{X}|^{2}|\mathcal{A}|T^{2}_{max}\right)\end{split}$ (27)
@@
$\begin{split}&(A.2)\\ &=\sum_{k,t,x,a}\sigma_{k,t}(x,a)\Big|\left<\left(P^{\tau}-\tilde{P}^{\tau}_{k}\right)(x,a),\left(V^{\tilde{P}^{\tau}_{k}}_{\pi_{k}}(\tilde{c}_{k})-V^{{P}^{\tau}}_{\pi_{k}}({c})\right)\right>\Big|\\ &\leq\underbrace{2\sqrt{L}\sum_{k,t,x,a}\frac{\sigma_{k,t}(x,a)}{\sqrt{N^{\tau}_{k}(x,a)\vee 1}}\left<\sqrt{\tilde{P}^{\tau}_{k}(x,a)},\Big|V^{\tilde{P}_{k}^{\tau}}_{\pi_{k}}(\tilde{c}_{k})-V^{{P}^{\tau}}_{\pi_{k}}({c})\Big|\right>}_{(B.1)}\\ &+\underbrace{\frac{14L}{3}\sum_{k,t,x,a}\sigma_{k,t}(x,a)\sum_{y\in\mathcal{X}}\frac{\Big|\left(V^{\tilde{P}_{k}^{\tau}}_{\pi_{k}}(\tilde{c}_{k},y)-V^{{P}^{\tau}}_{\pi_{k}}({c},y)\right)\Big|}{{N^{\tau}_{k}(x,a)\vee 1}}}_{(B.2)}\end{split}$ (28)
@@
$\begin{split}&(B.2)\\ &=\frac{14L}{3}\sum_{k,t,x,a}\sigma_{k,t}(x,a)\sum_{y\in\mathcal{X}}\frac{\Big|\left(V^{\tilde{P}_{k}^{\tau}}_{\pi_{k}}(\tilde{c}_{k},y)-V^{{P}^{\tau}}_{\pi_{k}}({c},y)\right)\Big|}{{N^{{\tau}}_{k}(x,a)\vee 1}}\\ &\overset{\overset{(a)}{<}}{\sim}|\mathcal{X}|\sum_{k,t,x,a}\frac{\sigma_{k,t}(x,a)}{{N^{{\tau}}_{k}(x,a)\vee 1}}\Big|\Big(V^{\tilde{P}_{k}^{\tau}}_{\pi_{k}}({c},y)+\frac{T_{max}}{p-p^{s}}\Theta_{k}^{\pi_{k}}(\tilde{P}^{\tau}_{k})\\ &+V^{{P}^{\tau}}_{\pi_{k}}({c},y)\Big)\Big|\\ &\overset{\overset{(b)}{<}}{\sim}|\mathcal{X}|\left(2T_{max}+\frac{T^{2}_{max}|\mathcal{X}|L}{p-p^{s}}\right)\sum_{k=1}^{K}\mathbb{E}^{x_{0}}_{\pi_{k}}\left[\sum_{t=0}^{T_{max}-1}\frac{1}{{N^{{\tau}}_{k}(x,a)\vee 1}};P^{\tau}\right]\\ &\overset{\overset{(c)}{<}}{\sim}|\mathcal{X}|\left(2T_{max}+\frac{T^{2}_{max}|\mathcal{X}|L}{p-p^{s}}\right)\tilde{\mathcal{O}}\left(|\mathcal{X}||\mathcal{A}|T_{max}\right)\\ &\overset{\overset{}{<}}{\sim}\tilde{\mathcal{O}}\left(\frac{T^{3}_{max}|\mathcal{X}|^{2}|\mathcal{A}|}{p-p^{s}}\right)\end{split}$ (29)
Relation $(a)$ is from the definition of $V^{\tilde{P}_{k}^{\tau}}_{\pi_{k}}(\tilde{c}_{k},x)$ and $\Theta_{k}^{\pi_{k}}(\tilde{P}^{\tau}_{k}):=\mathds{E}_{\pi_{k}}^{y}\left[\sum_{t=0}^{T_{max}-1}\hat{\epsilon}^{\tau}_{k}(X^{\tau}_{k,t},A^{\tau}_{k,t});\tilde{P}^{\tau}_{k}\right]$. Relation $(b)$ follows from the fact that $\Theta_{k}^{\pi_{k}}(\tilde{P}^{\tau}_{k})\overset{{<}}{\sim}T_{max}|\mathcal{X}|L$. Relation $(c)$ is due to Lemma 20.
@@
$\begin{split}&(B.1)\\ &=2\sqrt{L}\sum_{k,t,x,a}\frac{\sigma_{k,t}(x,a)}{\sqrt{N^{\tau}_{k}(x,a)\vee 1}}\left<\sqrt{\tilde{P}^{\tau}_{k}(x,a)},\Big|V^{\tilde{P}_{k}^{\tau}}_{\pi_{k}}(\tilde{c}_{k})-V^{{P}^{\tau}}_{\pi_{k}}({c})\Big|\right>\\ &\overset{\overset{(a)}{<}}{\sim}\sum_{k,t,x,a}\sigma_{k,t}(x,a)\frac{\sqrt{|\mathcal{X}|}\sqrt{\left<\tilde{P}^{{\tau}}_{k}(x,a),\left(V^{\tilde{P}_{k}^{\tau}}_{\pi_{k}}(\tilde{c}_{k})-V^{{P}^{\tau}}_{\pi_{k}}({c})\right)^{2}\right>}}{\sqrt{N^{{\tau}}_{k}(x,a)\vee 1}}\\ &\overset{\overset{(b)}{<}}{\sim}\sqrt{|\mathcal{X}|}\left(\sum_{t,k,x,a}\frac{\sigma_{k,t}(x,a)}{N^{{\tau}}_{k}(x,a)}\right)^{\frac{1}{2}}\\ &\times\left(\sum_{k,t,x,a,y}\sigma_{k,t}(x,a)\tilde{P}^{{\tau}}_{k}(x,a,y)\left(V^{\tilde{P}_{k}^{\tau}}_{\pi_{k}}(\tilde{c}_{k},y)-V^{{P}^{\tau}}_{\pi_{k}}({c},y)\right)^{2}\right)^{\frac{1}{2}}\\ &\overset{\overset{(c)}{<}}{\sim}\tilde{\mathcal{O}}\left(\sqrt{|\mathcal{X}|^{2}|\mathcal{A}|T_{max}}\right)\\ &\times\left(\sum_{k,t,y,a}\sigma_{k,t+1}(y,a)\left(V^{\tilde{P}_{k}^{\tau}}_{\pi_{k}}(\tilde{c}_{k},y)-V^{{P}^{\tau}}_{\pi_{k}}({c},y)\right)^{2}\right)^{\frac{1}{2}}\end{split}$ (30)
@@
$\begin{split}&\left(V^{\tilde{P}_{k}^{\tau}}_{\pi_{k}}(\tilde{c}_{k},y)-V^{{P}^{\tau}}_{\pi_{k}}({c},y)\right)^{2}\\ &\overset{\overset{(a)}{<}}{\sim}\tilde{\mathcal{O}}\left(\frac{1}{p-p^{s}}T^{2}_{max}|\mathcal{X}|L\right)\left(V^{{P}^{\tau}}_{\pi_{k}}({c},y)-V^{\tilde{P}_{k}^{\tau}}_{\pi_{k}}(\tilde{c}_{k},y)\right)\end{split}$
@@
$\begin{split}&X=\sum_{k,t,x,a}\sigma_{k,t}(x,a)\left|\left<\left(P^{\tau}-\tilde{P}^{\tau}_{k}\right)(x,a),\left(V^{\tilde{P}_{k}^{\tau}}_{\pi_{k}}(\tilde{c}_{k})-V^{{P}^{\tau}}_{\pi_{k}}({c})\right)\right>\right|,\\ &a=\frac{T^{3}_{max}|\mathcal{X}|^{2}|\mathcal{A}|}{p-p^{s}}\\ &+\sqrt{\frac{|\mathcal{X}|^{3}|\mathcal{A}|T^{4}_{max}L}{p-p^{s}}}\Big(\sum_{k}\Big(V^{P^{\tau}}_{\pi_{k}}(c,x_{0})-V^{\tilde{P}_{k}^{\tau}}_{\pi_{k}}(\tilde{c}_{k},x_{0})\Big)\Big)^{\frac{1}{2}},\\ &b=\sqrt{\frac{|\mathcal{X}|^{3}|\mathcal{A}|T^{4}_{max}L}{p-p^{s}}}\end{split}$
@@
$\begin{split}&(A.2)\\ &=\sum_{k,t,x,a}\sigma_{k,t}(x,a)\Big|\left<\left(P^{\tau}-\tilde{P}^{\tau}_{k}\right)(x,a),\left(V^{\tilde{P}_{k}^{\tau}}_{\pi_{k}}(\tilde{c}_{k})-V^{{P}^{\tau}}_{\pi_{k}}({c})\right)\right>\Big|\\ &\overset{\overset{(a)}{<}}{\sim}\frac{T^{3}_{max}|\mathcal{X}|^{2}|\mathcal{A}|}{p-p^{s}}\\ &+\sqrt{\frac{|\mathcal{X}|^{3}|\mathcal{A}|T^{4}_{max}L}{p-p^{s}}}\Big(\sum_{k}\Big(V^{P^{\tau}}_{\pi_{k}}(c,x_{0})-V^{\tilde{P}_{k}^{\tau}}_{\pi_{k}}(\tilde{c}_{k},x_{0})\Big)\Big)^{\frac{1}{2}}\\ &+\frac{|\mathcal{X}|^{3}|\mathcal{A}|T^{4}_{max}L}{p-p^{s}}\end{split}$ (33)
@@
$\begin{split}&(II)\\ &\overset{\overset{}{<}}{\sim}\sqrt{|\mathcal{X}|^{2}|\mathcal{A}|T^{3}_{max}K}+|\mathcal{X}|^{2}|\mathcal{A}|T^{2}_{max}+\frac{T^{3}_{max}|\mathcal{X}|^{2}|\mathcal{A}|}{p-p^{s}}\\ &+\sqrt{\frac{|\mathcal{X}|^{3}|\mathcal{A}|T^{4}_{max}L}{p-p^{s}}}\Big(\sum_{k}\Big(V^{P^{\tau}}_{\pi_{k}}(c,x_{0})-V^{\tilde{P}_{k}^{\tau}}_{\pi_{k}}(\tilde{c}_{k},x_{0})\Big)\Big)^{\frac{1}{2}}\\ &+\frac{|\mathcal{X}|^{3}|\mathcal{A}|T^{4}_{max}L}{p-p^{s}}\end{split}$ (34)

### v3

@@ 6 Conclusion and Future Works
$\begin{split}&R(K)\\ &=\sum_{k=1}^{K}\ \big(V^{P}_{\pi_{k}}(c,x_{0})-V^{P}_{\pi^{*}}(c,x_{0})\big)\\ &\overset{(a)}{=}\sum_{k=1}^{K}\ \big(V^{P^{\tau}}_{\pi_{k}}(c,x_{0})-V^{{P}^{\tau}}_{\pi^{*}}(c,x_{0})\big)\\ &=\sum_{k=1}^{K}\ \big(V^{P^{\tau}}_{\pi_{k}}(c,x_{0})-V^{\tilde{P}_{k}^{\tau}}_{\pi_{k}}(\tilde{c}_{k},x_{0})+V^{\tilde{P}_{k}^{\tau}}_{\pi_{k}}(\tilde{c}_{k},x_{0})-V^{P^{\tau}}_{\pi^{*}}(c,x_{0})\big)\\ &\overset{(b)}{\leq}\sum_{k=1}^{K}\ \big(V^{P^{\tau}}_{\pi_{k}}(c,x_{0})-V^{\tilde{P}_{k}^{\tau}}_{\pi_{k}}(\tilde{c}_{k},x_{0})\big)\\ &\overset{(c)}{\leq}\underbrace{\sum_{k=1}^{K}\ \mathbb{E}^{x_{0}}_{\pi_{k}}\sum_{t=0}^{T_{max}-1}\left[\left(c(X^{\tau}_{k,t},A^{\tau}_{k,t})-\tilde{c}_{k}(X^{\tau}_{k,t},A^{\tau}_{k,t})\right);P^{\tau}\right]}_{\text{(I)}}\\ &+\underbrace{\sum_{k=1}^{K}\mathbb{E}^{x_{0}}_{\pi_{k}}\sum_{t=0}^{T_{max}-1}\left[\left<\left|\big(P^{\tau}-\tilde{P}^{\tau}_{k}\big)(X^{\tau}_{k,t},A^{\tau}_{k,t})\right|,V^{\tilde{P}_{k}^{\tau}}_{\pi_{k}}(\tilde{c}_{k})\right>;P^{\tau}\right]}_{\text{(II)}}.\end{split}$ (25)
@@
We now bound the terms $(I)$ and $(II)$ within the good event $\mathcal{G}$, i.e., with probability at least $(1-4\delta)$.
We bound the term $(I)$ as follows.
$\begin{split}(I)&=\sum_{k=1}^{K}\ \mathbb{E}^{x_{0}}_{\pi_{k}}\sum_{t=0}^{T_{max}-1}\left[\left(c(X^{\tau}_{k,t},A^{\tau}_{k,t})-\tilde{c}_{k}(X^{\tau}_{k,t},A^{\tau}_{k,t})\right);P^{\tau}\right]\\ &\leq\frac{4T_{max}}{p-p^{s}}\sum_{k=1}^{K}\mathbb{E}^{x_{0}}_{\pi_{k}}\left[\sum_{t=0}^{T_{max}-1}\hat{\epsilon}^{\tau}_{k}(X^{\tau}_{k,t},A^{\tau}_{k,t});P^{\tau}\right]\\ &\overset{\overset{(a)}{<}}{\sim}\tilde{\mathcal{O}}\left(\frac{1}{p-p^{s}}\left(|\mathcal{X}|\sqrt{|\mathcal{A}|T^{3}_{max}K}+|\mathcal{X}|^{2}|\mathcal{A}|T^{2}_{max}\right)\right).\end{split}$ (26)
@@
$\begin{split}&(II)\\ &=\sum_{k,t}\mathbb{E}^{x_{0}}_{\pi_{k}}\left[\sum_{y\in\mathcal{X}}\big|\left(P^{\tau}-\tilde{P}^{\tau}_{k}\right)(X^{\tau}_{k,t},A^{\tau}_{k,t},y)\big|V^{\tilde{P}_{k}^{\tau}}_{\pi_{k}}(\tilde{c}_{k},y)\ ;P^{\tau}\right]\\ &\overset{}{=}\sum_{k,t,x,a}\sigma_{k,t}(x,a)\sum_{y\in\mathcal{X}}\big|\left(P^{\tau}-\tilde{P}^{\tau}_{k}\right)(x,a,y)\big|V^{\tilde{P}_{k}^{\tau}}_{\pi_{k}}(\tilde{c}_{k},y)\\ &=\underbrace{\sum_{k,t,x,a}\sigma_{k,t}(x,a)\sum_{y\in\mathcal{X}}\big|\left(P^{\tau}-\tilde{P}^{\tau}_{k}\right)(x,a,y)\big|V^{{P}^{\tau}}_{\pi_{k}}({c},y)}_{(A.1)}\\ &+\underbrace{\sum_{k,t,x,a}\sigma_{k,t}(x,a)\left<\big|\left(P^{\tau}-\tilde{P}^{\tau}_{k}\right)(x,a)\big|,\big(V^{\tilde{P}_{k}^{\tau}}_{\pi_{k}}(\tilde{c}_{k})-V^{{P}^{\tau}}_{\pi_{k}}({c})\big)\right>}_{(A.2)}.\end{split}$ (27)
The term $(A.1)$ can be bounded as:
$\begin{split}(A.1)&=\sum_{k,t,x,a}\sigma_{k,t}(x,a)\sum_{y\in\mathcal{X}}\big|\left(P^{\tau}-\tilde{P}^{\tau}_{k}\right)(x,a,y)\big|V^{{P}^{\tau}}_{\pi_{k}}({c},y)\\ &\overset{(a)}{\leq}T_{max}\sum_{k,t,x,a}\sigma_{k,t}(x,a)\sum_{y\in\mathcal{X}}\hat{\epsilon}^{\tau}_{k}(X^{\tau}_{k,t},A^{\tau}_{k,t},y)\\ &\overset{\overset{(b)}{<}}{\sim}\tilde{\mathcal{O}}\left(\sqrt{|\mathcal{X}|^{2}|\mathcal{A}|T^{3}_{max}K}+|\mathcal{X}|^{2}|\mathcal{A}|T^{2}_{max}\right).\end{split}$ (28)
@@
$\begin{split}&(A.2)\\ &=\sum_{k,t,x,a}\sigma_{k,t}(x,a)\Big|\left<\left(P^{\tau}-\tilde{P}^{\tau}_{k}\right)(x,a),\left(V^{\tilde{P}^{\tau}_{k}}_{\pi_{k}}(\tilde{c}_{k})-V^{{P}^{\tau}}_{\pi_{k}}({c})\right)\right>\Big|\\ &\leq\underbrace{2\sqrt{L}\sum_{k,t,x,a}\frac{\sigma_{k,t}(x,a)}{\sqrt{N^{\tau}_{k}(x,a)\vee 1}}\left<\sqrt{\tilde{P}^{\tau}_{k}(x,a)},\Big|V^{\tilde{P}_{k}^{\tau}}_{\pi_{k}}(\tilde{c}_{k})-V^{{P}^{\tau}}_{\pi_{k}}({c})\Big|\right>}_{(B.1)}\\ &+\underbrace{\frac{14L}{3}\sum_{k,t,x,a}\sigma_{k,t}(x,a)\sum_{y\in\mathcal{X}}\frac{\Big|\left(V^{\tilde{P}_{k}^{\tau}}_{\pi_{k}}(\tilde{c}_{k},y)-V^{{P}^{\tau}}_{\pi_{k}}({c},y)\right)\Big|}{{N^{\tau}_{k}(x,a)\vee 1}}}_{(B.2)}.\end{split}$ (29)
@@
$\begin{split}&(B.2)\\ &=\frac{14L}{3}\sum_{k,t,x,a}\sigma_{k,t}(x,a)\sum_{y\in\mathcal{X}}\frac{\Big|\left(V^{\tilde{P}_{k}^{\tau}}_{\pi_{k}}(\tilde{c}_{k},y)-V^{{P}^{\tau}}_{\pi_{k}}({c},y)\right)\Big|}{{N^{{\tau}}_{k}(x,a)\vee 1}}\\ &\overset{\overset{(a)}{<}}{\sim}\sum_{k,t,x,a}\frac{\sigma_{k,t}(x,a)}{{N^{{\tau}}_{k}(x,a)\vee 1}}\sum_{y\in\mathcal{X}}\Big|\Big(V^{\tilde{P}_{k}^{\tau}}_{\pi_{k}}({c},y)+\frac{T_{max}}{p-p^{s}}\Theta_{k}^{\pi_{k}}(\tilde{P}^{\tau}_{k})\\ &+V^{{P}^{\tau}}_{\pi_{k}}({c},y)\Big)\Big|\\ &\overset{\overset{(b)}{<}}{\sim}|\mathcal{X}|\left(2T_{max}+\frac{T^{2}_{max}|\mathcal{X}|L}{p-p^{s}}\right)\sum_{k=1}^{K}\mathbb{E}^{x_{0}}_{\pi_{k}}\left[\sum_{t=0}^{T_{max}-1}\frac{1}{{N^{{\tau}}_{k}(x,a)\vee 1}};P^{\tau}\right]\\ &\overset{\overset{(c)}{<}}{\sim}|\mathcal{X}|\left(2T_{max}+\frac{T^{2}_{max}|\mathcal{X}|L}{p-p^{s}}\right)\tilde{\mathcal{O}}\left(|\mathcal{X}||\mathcal{A}|T_{max}\right)\\ &\overset{\overset{}{<}}{\sim}\tilde{\mathcal{O}}\left(\frac{T^{3}_{max}|\mathcal{X}|^{2}|\mathcal{A}|}{p-p^{s}}\right).\end{split}$ (30)
Relation $(a)$ is from the definition of $V^{\tilde{P}_{k}^{\tau}}_{\pi_{k}}(\tilde{c}_{k},x)$ and $\Theta_{k}^{\pi_{k}}(\tilde{P}^{\tau}_{k}):=\mathds{E}_{\pi_{k}}^{y}\left[\sum_{t=0}^{T_{max}-1}\hat{\epsilon}^{\tau}_{k}(X^{\tau}_{k,t},A^{\tau}_{k,t});\tilde{P}^{\tau}_{k}\right]$. Relation $(b)$ follows from the fact that $V_{\pi}^{P}(c,x)\leq T_{max}$ for all $\pi,P$ and $\Theta_{k}^{\pi_{k}}(\tilde{P}^{\tau}_{k})\overset{{<}}{\sim}T_{max}|\mathcal{X}|L$. Relation $(c)$ is due to Lemma 20.
@@
$\begin{split}&(B.1)\\ &=2\sqrt{L}\sum_{k,t,x,a}\frac{\sigma_{k,t}(x,a)}{\sqrt{N^{\tau}_{k}(x,a)\vee 1}}\left<\sqrt{\tilde{P}^{\tau}_{k}(x,a)},\Big|V^{\tilde{P}_{k}^{\tau}}_{\pi_{k}}(\tilde{c}_{k})-V^{{P}^{\tau}}_{\pi_{k}}({c})\Big|\right>\\ &\overset{\overset{(a)}{<}}{\sim}\sum_{k,t,x,a}\sigma_{k,t}(x,a)\frac{\sqrt{|\mathcal{X}|}\sqrt{\left<\tilde{P}^{{\tau}}_{k}(x,a),\left(V^{\tilde{P}_{k}^{\tau}}_{\pi_{k}}(\tilde{c}_{k})-V^{{P}^{\tau}}_{\pi_{k}}({c})\right)^{2}\right>}}{\sqrt{N^{{\tau}}_{k}(x,a)\vee 1}}\\ &\overset{\overset{(b)}{<}}{\sim}\sqrt{|\mathcal{X}|}\left(\sum_{t,k,x,a}\frac{\sigma_{k,t}(x,a)}{N^{{\tau}}_{k}(x,a)}\right)^{\frac{1}{2}}\\ &\times\left(\sum_{k,t,x,a,y}\sigma_{k,t}(x,a)\tilde{P}^{{\tau}}_{k}(x,a,y)\left(V^{\tilde{P}_{k}^{\tau}}_{\pi_{k}}(\tilde{c}_{k},y)-V^{{P}^{\tau}}_{\pi_{k}}({c},y)\right)^{2}\right)^{\frac{1}{2}}\\ &\overset{\overset{(c)}{<}}{\sim}\tilde{\mathcal{O}}\left(\sqrt{|\mathcal{X}|^{2}|\mathcal{A}|T_{max}}\right)\\ &\times\left(\sum_{k,t,y,a}\sigma_{k,t+1}(y,a)\left(V^{\tilde{P}_{k}^{\tau}}_{\pi_{k}}(\tilde{c}_{k},y)-V^{{P}^{\tau}}_{\pi_{k}}({c},y)\right)^{2}\right)^{\frac{1}{2}}.\end{split}$ (31)
@@
$\begin{split}&\left(V^{\tilde{P}_{k}^{\tau}}_{\pi_{k}}(\tilde{c}_{k},y)-V^{{P}^{\tau}}_{\pi_{k}}({c},y)\right)^{2}\\ &\overset{\overset{(a)}{<}}{\sim}\tilde{\mathcal{O}}\left(\frac{1}{p-p^{s}}T^{2}_{max}|\mathcal{X}|L\right)\left(V^{{P}^{\tau}}_{\pi_{k}}({c},y)-V^{\tilde{P}_{k}^{\tau}}_{\pi_{k}}(\tilde{c}_{k},y)\right).\end{split}$
@@
$\begin{split}&X=\sum_{k,t,x,a}\sigma_{k,t}(x,a)\left|\left<\left(P^{\tau}-\tilde{P}^{\tau}_{k}\right)(x,a),\left(V^{\tilde{P}_{k}^{\tau}}_{\pi_{k}}(\tilde{c}_{k})-V^{{P}^{\tau}}_{\pi_{k}}({c})\right)\right>\right|,\\ &a=\frac{T^{3}_{max}|\mathcal{X}|^{2}|\mathcal{A}|}{p-p^{s}}\\ &+\sqrt{\frac{|\mathcal{X}|^{3}|\mathcal{A}|T^{4}_{max}L}{p-p^{s}}}\Big(\sum_{k}\Big(V^{P^{\tau}}_{\pi_{k}}(c,x_{0})-V^{\tilde{P}_{k}^{\tau}}_{\pi_{k}}(\tilde{c}_{k},x_{0})\Big)\Big)^{\frac{1}{2}},\\ &b=\sqrt{\frac{|\mathcal{X}|^{3}|\mathcal{A}|T^{4}_{max}L}{p-p^{s}}}.\end{split}$
@@
$\begin{split}&(A.2)\\ &=\sum_{k,t,x,a}\sigma_{k,t}(x,a)\Big|\left<\left(P^{\tau}-\tilde{P}^{\tau}_{k}\right)(x,a),\left(V^{\tilde{P}_{k}^{\tau}}_{\pi_{k}}(\tilde{c}_{k})-V^{{P}^{\tau}}_{\pi_{k}}({c})\right)\right>\Big|\\ &\overset{\overset{(a)}{<}}{\sim}\frac{T^{3}_{max}|\mathcal{X}|^{2}|\mathcal{A}|}{p-p^{s}}\\ &+\sqrt{\frac{|\mathcal{X}|^{3}|\mathcal{A}|T^{4}_{max}L}{p-p^{s}}}\Big(\sum_{k}\Big(V^{P^{\tau}}_{\pi_{k}}(c,x_{0})-V^{\tilde{P}_{k}^{\tau}}_{\pi_{k}}(\tilde{c}_{k},x_{0})\Big)\Big)^{\frac{1}{2}}\\ &+\frac{|\mathcal{X}|^{3}|\mathcal{A}|T^{4}_{max}L}{p-p^{s}}.\end{split}$ (34)
@@
$\begin{split}&(II)\\ &\overset{\overset{}{<}}{\sim}\sqrt{|\mathcal{X}|^{2}|\mathcal{A}|T^{3}_{max}K}+|\mathcal{X}|^{2}|\mathcal{A}|T^{2}_{max}+\frac{T^{3}_{max}|\mathcal{X}|^{2}|\mathcal{A}|}{p-p^{s}}\\ &+\sqrt{\frac{|\mathcal{X}|^{3}|\mathcal{A}|T^{4}_{max}L}{p-p^{s}}}\Big(\sum_{k}\Big(V^{P^{\tau}}_{\pi_{k}}(c,x_{0})-V^{\tilde{P}_{k}^{\tau}}_{\pi_{k}}(\tilde{c}_{k},x_{0})\Big)\Big)^{\frac{1}{2}}\\ &+\frac{|\mathcal{X}|^{3}|\mathcal{A}|T^{4}_{max}L}{p-p^{s}}.\end{split}$ (35)